How is the Graph of velocity vs displacement in free fall?

Grade:11

1 Answers

Susmita
5 years ago
The equation of force is
Ma=Mg
Or,a=g
Or,dv/dt=g
Or,\frac{dv}{dx}\frac{dx}{dt}=g
Or,\frac{dv}{dx}v=g
Or,vdv=gdx
Or,\int vdv=\int gdx
Or,\frac{v^2}{2}=gx+c
Or,v^2 =2(gx+c)
Or,v2=2g(x+c/g)
This is equation of a parabola with vertex at (-c/g,0) and focus at (g/2,0)